OpenStudy (anonymous):

Let A and B be 2x2 matrices. Determine the matrix of the operator T: M->AMB on the space \(F^{2x2}\) of 2x2 matrices, with respect to the basis \((e_{11}, e_{12}, e_{21}, e_{22})\) of \(F^{2x2}\).

OpenStudy (anonymous):

This problem seems easier than I'm making it, I'm just getting mixed up. I know B transpose is involved, but I don't think \(B^tA\) is right.

OpenStudy (turingtest):

\[e_{11}=\left[\begin{matrix}1&0\\0&0\end{matrix}\right]\]\[e_{12}=\left[\begin{matrix}0&1\\0&0\end{matrix}\right]\]etc.?

OpenStudy (anonymous):

Yes, those are definitely the basis matrices, but I'm still lost as to figuring out the matrix of the operator.

hero (hero):

I only know what I was taught in school, sorry.

OpenStudy (anonymous):

I'm so confused as to why this problem is hard. It seems simple. And it's question 1 of the section of the book >.<

hero (hero):

of which book?

hero (hero):

Title, Author, and Edition Please

OpenStudy (anonymous):

Artin's Algebra, second edition. Question 4.2.1

OpenStudy (asnaseer):

not sure I understand the question. Is it asking for the result of this multiplication?\[\left[\begin{matrix}a_{11}&a_{12}\\a_{21}&a_{22}\end{matrix}\right]\left[\begin{matrix}e_{11}&e_{12}\\e_{21}&e_{22}\end{matrix}\right]\left[\begin{matrix}b_{11}&b_{12}\\b_{21}&b_{22}\end{matrix}\right]\]

OpenStudy (anonymous):

Some background for the question. We're talking about linear operators, which are linear transformations from a vector space back to itself. It is a theorem that any linear operator can be represented as left multiplication by a matrix. The question is to find the matrix for this given operator.

hero (hero):

Can't find an online version of the book anywhere

OpenStudy (asnaseer):

but the operator just says take some matrix M and transform it by applying AMB - correct?

OpenStudy (anonymous):

asna: Correct. But we need to find a matrix for that operator that transforms it by left multiplication only. So, we need to find what matrix, call it P, such that PM=AMB. Now that I've worded it that way... maybe I can just manipulate that equivalence to figure it out? Will try it out.

OpenStudy (asnaseer):

ok - so yes - if you carry out the matrix multiplications then PM and AMB will give you 2x2 matrices. so you then just have to compare corresponding elements to work out what P should be.

OpenStudy (anonymous):

I tried that, but it gets really messy and I can't figure it out =/

OpenStudy (anonymous):

In AMB you end up with, for example, the 1,1 entry has all four bases in it. So I don't know how you make one matrix that will lead to that.

OpenStudy (anonymous):

If we view B as a column operation matrix, then we can say that B transpose would be an equivalent row operation matrix, which is why I was thinking that B transpose was involved. But, I'm not sure quite how.

OpenStudy (asnaseer):

one second, if we write this out as matrix manipulations, then we get: \[\begin{align} PM &= AMB\\ P^{-1}PM&=P^{-1}AMB\\ M&=P^{-1}AMB\\ \end{align}\]which implies:\[P=A^{-1}\]and\[B=I\text{ (the identity matrix)}\]

OpenStudy (asnaseer):

sorry I meant P=A

OpenStudy (anonymous):

I don't think that it implies that. There are other ways for the equivalence to hold.

OpenStudy (asnaseer):

ok - let me try the messy algebraic way...

OpenStudy (anonymous):

Proof that that implication doesn't hold: Look at conjugation by an elementary matrix \(E\). We can have \(A=E^{-1}AE\) without it being the case that \(E=I\). For example, any permutation matrix.

OpenStudy (asnaseer):

true

OpenStudy (asnaseer):

hmmm... algebraic method is not leading anywhere sensible - there must be something obvious we are missing in the question.

OpenStudy (anonymous):

Right?? That's what I keep thinking, but I don't see it.

OpenStudy (asnaseer):

let me mull over for a while...